# softMathematical software

### v1

@@ Abstract.
The central concept of this paper is that of self-adhesivity defined in algebraic terms, which is a combinatorial reflection of the self-adhesivity concept studied earlier in context of polymatroids and information theory. The generalization also leads to a self-adhesivity operator defined on the hyper-level of CI frames. We answer some of the questions related to this approach and raise other open questions.
@@ 1. Introduction
This paper is devoted to the concept of self-adhesivity in the context of abstract conditional independence (CI) structures, which generalize/approximate classic probabilistic CI structures [29]. The concepts of adhesivity and self-adhesivity have been introduced by Matúš [23] in the context of polymatroids, which are special set functions playing a crucial role in combinatorial optimization [11]. Matúš’s motivation came from information theory [35], more specifically, from the area of information-theoretical inequalities. The self-adhesivity concept for polymatroids is a kind of abstraction of the method used to derive the first non-Shannon information inequality [37], which method later became known in information-theoretical community as the so-called copy lemma [10]. The method leads to the application of tools of polyhedral geometry (and linear programming solvers) to derive new inequalities [8].
In this paper we introduce the concept of self-adhesivity for abstract CI structures, which is a combinatorial reflection of the self-adhesivity concept for polymatroids. The idea comes from [2] and has already been applied in the context of Gaussian CI structures [3]. Nonetheless, our paper puts even more emphasis on the abstract view on (the families of) CI models and utilizes the results of lattice theory [1] for this purpose. More specifically, every (finite) lattice can alternatively be described in terms of a closure operation, or in terms of (abstract) functional dependence relation [19]. Thus, one can utilize the results from the theory of formal concept analysis [13]. This leads to an analogy with mathematical logic and to the effort to describe families of CI models in an “axiomatic way”, using implications among CI statements.
We introduce the notion of an abstract frame for CI structures and give several specific examples of such CI frames. We also introduce relevant algebraic operations with respect to which these abstract CI frames should be closed: these are copying and marginalization. From the point of view of our paper, the substantial algebraic operation is the set-theoretical intersection, which ensures that, for a fixed (finite) set of variables, the family of respective CI models forms a finite lattice. Then we introduce the concept of self-adhesivity, which is relative to a CI frame. The self-adhesivity concept induces special self-adhesivity operator of the hyper-level of CI frames. This operator assigns a new CI frame to a starting CI frame and this new frame is a shrinkage of the original frame. In case the starting CI frame is an outer approximation of the classic probabilistic CI frame, the application of the self-adhesivity operator yields an even tighter outer approximation of the probabilistic CI frame.
@@
Our approach also allows one to derive such CI implications directly, by means of performing the computations on the combinatorial level of CI frames. To contrast with Matúš’s original concept of self-adhesivity for polymatroids, note that the dimension of the polymatroidal cone grows exponentially in $n$ and the number of CI statements grows even a little bit faster. Still, our computational methods based on SAT solvers can deal with this increase much better than the linear programming solvers could which are used in the polymatroid setup.
In the paper, a specific example of the application of our combinatorial method is given. We applied the method to the question of identification of those extreme rays of the polymatroidal cone over $5$ variables which are not entropic (= do not have probabilistic representation). It appears that most of 1319 permutational types of these extreme rays lie outside the entropic region. More specifically, it follows from our computations that at most 154 of types of such extreme rays may be entropic. The proof is entirely computational and takes only 9 minutes to execute on an average laptop. Throughout the paper we also formulate remaining open mathematical problems related to our approach.
The structure of the paper is as follows. In Section 2, we define our concept of an abstract CI frame together with corresponding algebraic operations applied to CI models. We also give three standard examples of such CI frames and recall those particular concepts from lattice theory on which our presentation is based. The concept of self-adhesion relative to such an abstract CI frame is then introduced in Section 3, together with related basic observations and two other relevant algebraic operations applied to CI models. Section 4 is then devoted to the self-adhesivity operator and the idea of its iterating. The results of our computational experiments are presented in Section 5 and their application in information-theoretical context is explained in Section 6. Concluding remarks are presented in Section 7. More detailed information about our computations, source code and resulting data files can be found on the website:
https://cinet.link/data/selfadhe-lattices
@@ 2. Basic concepts
2.1. Abstract families of CI models
A standard elementary CI statement over $N$ is the statement claiming that, for some $K\subseteq N$ and distinct $i,j\in N\setminus K$, the variable $i$ is conditionally independent of the variable $j$ given $K$, traditionally denoted by $i\perp\!\!\!\perp j\,|\,K$. A converse conditional dependence statement (= negation) is the claim that $i$ and $j$ are conditionally dependent given $K$, traditionally denoted by $i\not\perp\!\!\!\perp j\,|\,K$. Thus, a triplet $\langle i,j|K\rangle$ of pairwise disjoint subsets of $N$ can be interpreted either as an independence statement or as a dependence statement.
@@ 2.1. Abstract families of CI models
By a (CI) frame we will understand a mapping which assigns a family over $N$ to every finite (variable) set $N$. Frames will be denoted using the fractal font without an argument: mapping $\mathfrak{f}:\,N\mapsto\mathfrak{f}(N)\in{\mathcal{P}}({\mathcal{P}}({\mathcal{K}}(N)))$.
Given two CI frames $\mathfrak{f}_{1}$ and $\mathfrak{f}_{2}$, the record $\mathfrak{f}_{1}\subseteq\mathfrak{f}_{2}$ means that $\mathfrak{f}_{1}(N)\subseteq\mathfrak{f}_{2}(N)$ for any (finite) variable set $N$.
The reader can recognize a clear hierarchy: a (CI) statement is an element of ${\mathcal{K}}(N)$, a model is an element of the power set of ${\mathcal{K}}(N)$ and a family is an element of the power set of the power set of ${\mathcal{K}}(N)$ (= the double power set). [^1] A frame corresponds to what was called a “property” in the PhD thesis [2, § 2.2].
@@
The next step is to introduce some “algebraic” operations with abstract CI models. The goal is to study various frames which are closed under such operations. Later, in Section 2.2, we give specific examples of abstract CI frames we have in mind.
Definition 2.2 (copying, marginalization, intersection, duality).
  • Any bijection $\varphi:N\to M$ between two variable sets $N$ and $M$ can naturally be extended to a bijection between their power sets and to a bijection between ${\mathcal{K}}(N)$ and ${\mathcal{K}}(M)$: put $\varphi(ij|K)\,:=\,\varphi(ij)|\varphi(K)$ for any $ij|K\in{\mathcal{K}}(N)$.
  • Any model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ over $N$ is then transformed by $\varphi$ to a model over $M$: put $\varphi({\mathcal{M}}):=\{\,\varphi(ij|K)\,:\ ij|K\in{\mathcal{M}}\,\}\subseteq{\mathcal{K}}(M)$. This is the $\varphi$-copy of ${\mathcal{M}}$. We say that a CI frame $\mathfrak{f}$ is closed under copying if, for any model ${\mathcal{M}}\in\mathfrak{f}(N)$ and any bijection $\varphi:N\to M$, one has $\varphi({\mathcal{M}})\in\mathfrak{f}(M)$.
  • Given a model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ over $N$ and  $M\subseteq N$, its marginal for $M$, denoted by ${\mathcal{M}}^{\downarrow M}$, is simply its restriction to ${\mathcal{K}}(M)$: put ${\mathcal{M}}^{\downarrow M}:={\mathcal{M}}\cap{\mathcal{K}}(M)$. A CI frame $\mathfrak{f}$ is closed under marginalization if, for any model ${\mathcal{M}}\in\mathfrak{f}(N)$ and any subset $M\subseteq N$, one has ${\mathcal{M}}^{\downarrow M}\in\mathfrak{f}(M)$.
  • Given two models ${\mathcal{M}},{\mathcal{R}}\subseteq{\mathcal{K}}(N)$ over the same set $N$, their (set-theoretical) intersection is the model ${\mathcal{M}}\cap{\mathcal{R}}$, again a model over $N$. A CI frame $\mathfrak{f}$ will be called closed under intersection if, for any two models ${\mathcal{M}},{\mathcal{R}}\in\mathfrak{f}(N)$, one has ${\mathcal{M}}\cap{\mathcal{R}}\in\mathfrak{f}(N)$.
  • Given a variable set $N$, a duality mapping from ${\mathcal{K}}(N)$ to ${\mathcal{K}}(N)$ is as follows: $ij|K\in{\mathcal{K}}(N)$ is assigned the pair $(ij|K)^{\star}:=ij|L$, where $L\,:=\,N\setminus(K\cup ij)$. Given a model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$, the dual model ${\mathcal{M}}^{\star}$ is its image by this mapping. A CI frame $\mathfrak{f}$ is closed under duality if ${\mathcal{M}}\in\mathfrak{f}(N)$ implies ${\mathcal{M}}^{\star}\in\mathfrak{f}(N)$.
Since ${\mathcal{K}}(N)=\emptyset$ if $|N|\leq 1$, the concepts of copying and duality make real sense only in case $|N|\geq 2$. On the other hand, to introduce later concepts from Section 3 one needs to define marginalization to sets $M$ with $|M|\leq 1$.
Two basic “algebraic” operations with abstract CI models are copying and marginalization. All frames we consider in this paper are closed under these two operations. Most of the considered CI frames are also closed under intersection, which implies that the respective families of models are lattices (see Section 2.3). A more complicated adhesivity concept will be discussed later (Section 3).
2.2. Examples of abstract CI frames
There are several distinct frames we consider in this paper.
2.2.1. Discrete probabilistic CI models
A prototypical example of a CI frame is the frame of discrete probabilistic CI structures. Here we define them specifically.
@@ 2.2.1. Discrete probabilistic CI models
This defines the discrete probabilistic CI frame $\mathfrak{pr}$.
@@
$\mathfrak{pr}$ is closed under copying, marginalization, and intersection.
Proof 2.4.
@@
On the other hand, the CI frame $\mathfrak{pr}$ is not closed under duality. Recall a counter-example from [21, end of § 5].
@@
of values for $(x_{a},x_{b},x_{c},x_{d})$ implies the former fact. The latter fact follows from the property (I:13) in [31, § V.B] saying that the implication

### v3

@@ Abstract.
The central concept of this paper is that of self-adhesivity defined in algebraic terms, which is a combinatorial reflection of the self-adhesivity concept studied earlier in context of polymatroids and information theory. The generalization also leads to a self-adhesivity operator defined on the meta-level of CI frames. We answer some of the questions related to this approach and raise other open questions.
@@ 1. Introduction
This paper is devoted to the concept of self-adhesivity in the context of abstract conditional independence (CI) structures, which generalize/approximate classic probabilistic CI structures [37]. The concepts of adhesivity and self-adhesivity have been introduced by Matúš [29] in the context of polymatroids, which are special set functions playing a crucial role in combinatorial optimization [14]. Matúš’s motivation came from information theory [41], more specifically, from the area of information-theoretical inequalities. The self-adhesivity concept for polymatroids is a kind of abstraction of the method used to derive the first non-Shannon information inequality [43], which method later became known in information-theoretical community as the so-called copy lemma [12]. The method led to the application of tools of polyhedral geometry (and linear programming solvers) to derive new inequalities [9].
In this paper we introduce the concept of self-adhesivity for abstract CI structures, which is a combinatorial reflection of the self-adhesivity concept for polymatroids. The idea comes from [5] and has already been applied in the context of Gaussian CI structures [6]. Nonetheless, our paper puts even more emphasis on the abstract view on (the families of) CI models and utilizes the results of lattice theory [4] for this purpose. More specifically, every finite lattice can alternatively be described in terms of a closure operation, or in terms of (abstract) functional dependence relation [24]. Thus, one can utilize the results from the theory of formal concept analysis [18]. This leads to an analogy with mathematical logic and to the effort to describe families of CI models in an axiomatic way, using implications among CI statements.
We introduce the notion of an abstract frame for CI structures and give several specific examples of such CI frames. We also introduce relevant algebraic operations with respect to which these abstract CI frames should be closed: these are copying and marginalization. From the point of view of our paper, the substantial algebraic operation is the set-theoretical intersection, which ensures that, for a fixed (finite) set of variables, the family of respective CI models forms a finite lattice. Then we introduce the concept of self-adhesivity, which is relative to a CI frame. The self-adhesivity concept induces special self-adhesivity operator of the meta-level of CI frames. This operator assigns a new CI frame to a starting CI frame and this new frame is a shrinkage of the original frame. In case the starting CI frame is an outer approximation of the classic probabilistic CI frame, the application of the self-adhesivity operator yields an even tighter outer approximation of the probabilistic CI frame.
@@
Our approach also allows one to derive such CI implications directly, by means of performing the computations on the combinatorial level of CI frames. To contrast with Matúš’s original concept of self-adhesivity for polymatroids, note that the dimension of the polymatroidal cone grows exponentially in $n$ and the number of CI statements grows even faster. Still, our computational methods based on SAT solvers can deal with this increase much better than the linear programming solvers used in the polymatroid setup.
In the paper, a specific example of the application of our combinatorial method is given. We applied the method to the question of identification of those extreme rays of the polymatroidal cone over $5$ variables which are not entropic, which means that they do not have probabilistic representation. It appears that most of the 1319 permutational types of these extreme rays lie outside the entropic region. More specifically, it follows from our computations that at most 154 of types of such extreme rays may be entropic. The proof is entirely computational and takes about 330 seconds to execute on the author’s laptop. Throughout the paper we also formulate remaining open mathematical questions related to our approach.
The structure of the paper is as follows. In Section 2, we define our concept of an abstract CI frame together with basic algebraic operations applied to CI models. We also recall those particular concepts and results from lattice theory on which our algorithms are based and which allow one to introduce frame-based closures of CI models. In Section 3 we present examples of such abstract CI frames. The concept of self-adhesion relative to such an abstract CI frame is introduced in Section 4, together with related basic observations and two other relevant algebraic operations applied to CI models. Section 5 is devoted to the self-adhesivity operator and the idea of its iterating. In Section 6 we review our tools for computations. The results of our computational experiments are presented in Section 7 and their application in information-theoretical context is explained in Section 8. Concluding remarks are presented in Section 9. More detailed information about our computations, source code and resulting data files can be found on the website:
https://cinet.link/data/selfadhe-lattices  .
@@ 2. Basic concepts
2.1. Abstract families of CI models and the probabilistic CI frame
An elementary CI statement over a finite variable set $N$ is the statement claiming that, for some $K\subseteq N$ and distinct $i,j\in N\setminus K$, the variable $i$ is conditionally independent of the variable $j$ given $K$, traditionally denoted by $i\perp\!\!\!\perp j\,|\,K$. A converse conditional dependence statement (= negation) is the claim that $i$ and $j$ are conditionally dependent given $K$, traditionally denoted by $i\not\perp\!\!\!\perp j\,|\,K$. Thus, a triplet $\langle i,j|K\rangle$ of pairwise disjoint subsets of $N$ can be interpreted either as an independence statement or as a dependence statement.
@@ 2.1. Abstract families of CI models and the probabilistic CI frame
Note that this way of introducing CI statements implicitly means that any CI statement over a variable set $N$ is automatically also a CI statement over any variable set $M$ with $N\subseteq M$.
@@
By a (CI) frame we will understand a mapping which assigns a (CI) family over $N$ to each finite (variable) set $N$. Frames will be denoted using the fractal font without an argument: mapping $\mathfrak{f}:\,N\mapsto\mathfrak{f}(N)\in{\mathcal{P}}({\mathcal{P}}({\mathcal{K}}(N)))$. Given two CI frames $\mathfrak{f}_{1}$ and $\mathfrak{f}_{2}$, we write $\mathfrak{f}_{1}\subseteq\mathfrak{f}_{2}$ to mean $\mathfrak{f}_{1}(N)\subseteq\mathfrak{f}_{2}(N)$ for any (finite) variable set $N$.
The reader can recognize a clear hierarchy: a (CI) statement is an element of ${\mathcal{K}}(N)$, a model is an element of the power set of ${\mathcal{K}}(N)$ and a family is an element of the power set of the power set of ${\mathcal{K}}(N)$. [^1] A frame corresponds to what was called a “property” in the PhD thesis [5, § 2.2].
@@
The definition of a CI model is relative to a variable set $N$ and, strictly speaking, the model is a pair $(N,{\mathcal{M}})$ with ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$. For example, the empty set might be considered as a CI model over $\{a,b\}$ or over $\{a,b,c\}$. To simplify notation, we will usually rely on context to identify the variable set or mention it explicitly by means of the “frame notation” ${\mathcal{M}}\in\mathfrak{f}(N)$. If two variable sets are in inclusion $N\subseteq M$, then any CI model ${\mathcal{M}}$ over $N$ may be regarded as a model over $M$ as well. In the context of a frame, this ability to transition to larger variable sets is formalized by the operation of ascetic extension defined below.
Example 2.2.
Put $N:=\{a,b,c,d\}$ and :=${\mathcal{M}}:=\{\,ab|\emptyset,bc|\emptyset\,\}\subseteq{\mathcal{K}}(N)$. Then $N^{{\mathcal{M}}}=\{a,b,c\}$ and ${\mathcal{M}}$ is a CI model over this strict subset of $N$. Nonetheless, ${\mathcal{M}}$ is not a CI model over $\{a,b\}$ nor over $\{b,c,e\}$, where $e$ is a variable outside $N$. On the other hand, ${\mathcal{M}}$ is a model over any variable set $M$ with $\{a,b,c\}\subseteq M$.
@@
Observe that ${\mathcal{K}}(N)$ belongs to $\mathfrak{pr}(N)$ because it is induced by a random vector with completely independent components. This defines the discrete probabilistic CI frame $\mathfrak{pr}$.
It will be convenient to deal with CI statements with more variables in place of the singletons $i$ and $j$. Although the treatment of (probabilistic) CI statements can be extended to this more general case, in this paper the notation $I\perp\!\!\!\perp J|K$ for disjoint subsets of $N$ will be considered only as an abbreviation.
Definition 2.3 (global CI statement).
We interpret this as a global CI statement encoded within the CI model ${\mathcal{M}}$. In our examples and catalogues we also use a shorthand $[I,J|K]$ to denote the list of (elementary) CI statements described in (1).
Note that, in the context of the probabilistic frame $\mathfrak{pr}$, the abbreviating condition $I\perp\!\!\!\perp J\,|\,K\,\,[{\mathcal{M}}_{\mbox{\scriptsize\boldmath$\xi$}}]$ is equivalent to the requirement that $p_{IJK}(x)\cdot p_{K}(x)=p_{IK}(x)\cdot p_{JK}(x)$ holds for every $x\in\mbox{\sf X}_{N}$, which requirement is interpreted as the (global) probabilistic CI statement $I\perp\!\!\!\perp J\,|\,K\,\,[\mbox{\boldmath$\xi$}]$; see [37, Lemma 2.2].
The next step is to introduce five algebraic operations on abstract CI models. The goal is to study various CI frames which are closed under such operations.
Definition 2.4 (copying, marginalization, intersection, duality, ascetic extension).
  • Any bijection $\varphi:N\to M$ between two variable sets $N$ and $M$ can naturally be extended to a bijection between their power sets and to a bijection between ${\mathcal{K}}(N)$ and ${\mathcal{K}}(M)$: put $\varphi(ij|K)\,:=\,\varphi(ij)|\varphi(K)$ for every $ij|K\in{\mathcal{K}}(N)$.
  • Any model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ over $N$ is then transformed by $\varphi$ to a model over $M$: put
    $\varphi({\mathcal{M}}):=\{\,\varphi(ij|K)\,:\ ij|K\in{\mathcal{M}}\,\}\subseteq{\mathcal{K}}(M)\,.$
    This is the $\varphi$-copy of ${\mathcal{M}}$. We say that a CI frame $\mathfrak{f}$ is closed under copying if, for every model ${\mathcal{M}}\in\mathfrak{f}(N)$ and every bijection $\varphi:N\to M$, one has $\varphi({\mathcal{M}})\in\mathfrak{f}(M)$.
  • Given a model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ over $N$ and  $M\subseteq N$, its marginal for $M$, denoted by ${\mathcal{M}}^{\downarrow M}$, is simply its restriction to ${\mathcal{K}}(M)$: put ${\mathcal{M}}^{\downarrow M}:={\mathcal{M}}\cap{\mathcal{K}}(M)$. A CI frame $\mathfrak{f}$ is closed under marginalization if, for every model ${\mathcal{M}}\in\mathfrak{f}(N)$ and every subset $M\subseteq N$, one has ${\mathcal{M}}^{\downarrow M}\in\mathfrak{f}(M)$.
  • Given two models ${\mathcal{M}},{\mathcal{R}}\subseteq{\mathcal{K}}(N)$ over the same set $N$, their (set-theoretical) intersection is the model ${\mathcal{M}}\cap{\mathcal{R}}$, again a model over $N$. A CI frame $\mathfrak{f}$ will be called closed under intersection if, for every two models ${\mathcal{M}},{\mathcal{R}}\in\mathfrak{f}(N)$, one has ${\mathcal{M}}\cap{\mathcal{R}}\in\mathfrak{f}(N)$.
  • Given a variable set $N$, the duality map from ${\mathcal{K}}(N)$ to ${\mathcal{K}}(N)$ is as follows: $ij|K\in{\mathcal{K}}(N)$ is assigned the pair $(ij|K)^{\star}:=ij|L$, where $L\,:=\,N\setminus(K\cup ij)$. Given a model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$, the dual model ${\mathcal{M}}^{\star}$ is its image by this mapping. A CI frame $\mathfrak{f}$ is closed under duality if ${\mathcal{M}}\in\mathfrak{f}(N)$ implies ${\mathcal{M}}^{\star}\in\mathfrak{f}(N)$.
  • Given a model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ over $N$ and $N\subseteq M$, an extension of ${\mathcal{M}}$ to $M$ is any model ${\mathcal{R}}\subseteq{\mathcal{K}}(M)$ over $M$ such that ${\mathcal{R}}^{\downarrow N}={\mathcal{M}}$. A CI frame $\mathfrak{f}$ is closed under extension if, for every model ${\mathcal{M}}\in\mathfrak{f}(N)$ and every $M$ with $M\supseteq N$, there exists at least one extension ${\mathcal{R}}$ of ${\mathcal{M}}$ to $M$ with ${\mathcal{R}}\in\mathfrak{f}(M)$.
  • Given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and $N\subseteq M$, an extension ${\mathcal{R}}\subseteq{\mathcal{K}}(M)$ of ${\mathcal{M}}$ to $M$ will be called ascetic if ${\mathcal{R}}\setminus{\mathcal{M}}=\emptyset$. A CI frame $\mathfrak{f}$ is closed under ascetic extension if, for every model ${\mathcal{M}}\in\mathfrak{f}(N)$ and every variable set $M$ with $M\supseteq N$, one has ${\mathcal{M}}\in\mathfrak{f}(M)$, that is, the ascetic extension of ${\mathcal{M}}$ to $M$ belongs to $\mathfrak{f}(M)$.
Since ${\mathcal{K}}(N)=\emptyset$ if $|N|\leq 1$, the concepts of copying and duality make real sense only in case $|N|\geq 2$. On the other hand, to introduce self-adhesion in Section 4 one needs marginalization to sets $M$ with $|M|\leq 1$.
Note that if $N\subseteq M$ then every model ${\mathcal{M}}$ over $N$ has an ascetic extension to $M$, namely ${\mathcal{M}}$ itself, viewed as a CI model over $M$. Clearly, this is the least possible extension of ${\mathcal{M}}$ to $M$ and, thus, unique. The word “ascetic" was chosen because, in comparison with other extending algebraic operations, it adds nothing to ${\mathcal{M}}$. Clearly, a CI frame $\mathfrak{f}$ is closed under ascetic extension iff $N\subseteq M$ implies $\mathfrak{f}(N)\subseteq\mathfrak{f}(M)$. We leave it to the reader as an easy exercise to show that, if a CI frame $\mathfrak{f}$ is closed under extension and intersection, then it is closed under ascetic extension iff $N\subseteq M$ implies ${\mathcal{K}}(N)\in\mathfrak{f}(M)$.
Two basic algebraic operations with abstract CI models are copying and marginalization. Most of the considered CI frames in this paper are also closed under intersection, which implies that the respective families of models are lattices and one can introduce frame-based closures for CI models (see Section 2.3).
Our prototypical example of a CI frame is closed under all but one algebraic operations considered.
@@
The frame $\mathfrak{pr}$ is closed under copying, marginalization, intersection, and ascetic extension.
Proof.
@@
[…]
On the other hand, the CI frame $\mathfrak{pr}$ is not closed under duality. Recall a counter-example by Matúš from [26, end of § 5].
@@
of values for $(x_{a},x_{b},x_{c},x_{d})$ implies the former fact. The latter fact follows from the property (I:13) in [39, § V.B] saying that the implication